You have just received an advertisement from Corleone Inc., a “paycheck loan” service. Corleone will charge you a fee of 5% for a two-week loan (i.e. if you borrow $100, you must repay $105 in two weeks’ time). Assuming a 52-week year, what is the Effective Annual Rate (EAR) that Corleone charges (rounded to the nearest whole percent)
Answer #1

Ans Effective annual rate ( EAR) = ( 1+ interest rate )52 weeks / 2 weeks - 1

= ( 1+ 5 % ) 52/2 - 1

= ( 1+5%)26 - 1

= 3.555673 - 1 (1.05)26 = 3.555673

= 2.555673

= 255.56 %